# A Lanczos-Based Algorithmic Approach for Spike Detection in Large Sample Covariance Matrices

### v1

@@ 1. Introduction
Large sample covariance matrices play a fundamental role in high-dimensional data analysis. A widely studied and sophisticated model in the literature is the spiked covariance matrix model [13, 27], in which a finite number of spikes—eigenvalues that separate from the bulk of the spectrum—are introduced into the population covariance matrix $\Sigma$. In this model, a key task is to estimate the number of spikes using the associated sample covariance matrices. Such a problem has been studied in various settings; see Section 1.3 for a detailed review. However, all existing works leverage the spectral properties of $W$, relying on specific statistics constructed from its eigenvalues—either the prominent outlier eigenvalues or the relatively small bulk eigenvalues. Despite the potential statistical consistency of these approaches under certain conditions, they can be computationally intensive and prone to numerical errors due to the direct calculation of eigenvalues [^1] Modern eigensolvers are very accurate for symmetric matrices. The errors here stem from using individual eigevalues that may fluctuate wildly, while averages of eigenvalues have much smaller variance. and associated Monte Carlo estimations.
@@
In this paper, instead of relying on eigenvalue-based statistics, we introduce a novel algorithmic approach built upon the well-known Lanczos iteration [^2] Due to instabilities in the iteration we use reorthogonalization [39]. [34], which operates directly on the sample covariance matrix $W$ using random vectors, uniformly distributed on the hypersphere. Our proposed methods are both statistically consistent and numerically cheap while still being robust. Furthermore, as a byproduct of determining the number of spikes, we also obtain estimates for the asymptotic density of the empirical spectral distribution of $W$ and the asymptotic locations of the outliers in $W$.
@@
Our approach is based on the observation that, under mild assumptions, the number of spikes is, with high probability, identical to the number of poles of a complex function—specifically, the Stieltjes transform of a measure that connects the empirical spectral distribution (ESD) and its asymptotic counterpart, the asymptotic spectral distribution (ASD). This stochastic measure, which we refer to as the spiked ASD (see (1.11) and (5.1) below), preserves a scaled version of the continuous density of the ASD while accounting for the random outliers in the ESD. Consequently, a key challenge is to develop an efficient and robust estimator for the spiked ASD. Rather than directly relying on the eigenvalues of $W$, the central innovation of our method is the use of Cholesky factorization for the Jacobi matrix obtained by applying the Lanczos iteration with $W$ to several test vectors. This approach is motivated by the fact that the Stieltjes transform of the spiked ASD of $W$ can be approximately characterized by a new fixed-point equation, based on continued fraction theory, whose coefficients are derived from the entries of the Cholesky factorization of the Jacobi matrix associated with the spiked ASD (see Section 3.2). This fixed-point equation differs from the commonly-used self-consistent equation in random matrix theory [3, 30]. Crucially, this equation, along with the Lanczos algorithm, enables a numerical estimation of the Stieltjes transform of the ASD through an iterative process without directly computing the eigenvalues of $W$ (see Algorithm SR.2 and Section 3.2).
@@
On the techinical side, our proofs rely on three main ideas. The first one is the fact that the orthogonal polynomials for a generic limiting random matrix spectral distribution exhibit exponentially small error terms in their asymptotics [32, 52, 15]. These results are based on the analysis of the Fokas–Its–Kitaev Riemann–Hilbert characterization of orthogonal polynomials [26]. The second is the so-called anisotropic local laws for covariance matrices [30, 43] and their extension to spiked models [17]. The last technique is a recently developed perturbation theory [15] for orthogonal polynomials, which builds on the Fokas–Its–Kitaev Riemann–Hilbert theory and is particularly compatible with the local laws from random matrix theory.
@@
Figure 1. Left: The ESD of $\Sigma$, where $\Sigma$ is an $N\times N$ diagonal matrix with $N=6000$ and entries given by the quantiles of the density $\frac{1}{K}\frac{x^{4}+1}{x^{2}}\sqrt{x-0.1}\sqrt{4-x}$ with $K$ being a normalizing constant. The first three diagonal entries are modified to 7, 6, and 6, forming the spikes of $\Sigma$. Right: The ESD of $W$ defined in (1.1), with $c_{N}=0.1$, $M=60000$, and $X$ having iid standard normal entries. The ESD is compared against the estimated outliers, their locations, and the approximate ASD obtained using Algorithms P.2 and P.3 with parameters $k=200$ and $C=1$.
@@ 1.1. The model and problem formulation
where $\Sigma$ is a positive-definite deterministic matrix, referred to as the population covariance matrix, and $X$ is an $N\times M$ random matrix with centered, independent, and identically distributed (iid) entries. In this paper, we are interested in the high-dimensional setting, where $M$ is comparably large to $N$. Specifically, we assume the existence of a small constant $10<\tau<$ 1 such that the aspect ratio $c_{N}:=N/M$ satisfies
@@
We adopt the spiked covariance matrix model following the framework established in [14]. Let the spiked covariance matrix $\Sigma$ admit the following spectral decomposition
@@
We differentiate $\Sigma_{0}$ from $\Sigma$ because if a limit is desired for certain spectral statistics of (1.1), then additional assumptions will need to be imposed on $\Sigma_{0}$; see Assumption 2 for more details. As a historical note, we recall that if $\sigma_{i}\equiv 1$ for $1\leq i\leq N$, it is well known that the eigenvalues of $W_{0}$ obey the Marchenko-Pastur law [35]. For general $\Sigma_{0}$, they are governed by the so-called deformed Marchenko-Pastur law [3, 30], which, in turn, determines $\mu_{\mathrm{ASD}}$.
@@ 1.2. An overview of our approach
The entries of the semi-infinite Jacobi matrix $\mathcal{J}(\widehat{\mu}_{\mathbf{b}})$ associated with $\widehat{\mu}_{\mathbf{b}}$ (see (3.4)) and its Cholesky factor $\mathcal{L}(\widehat{\mu}_{\mathbf{b}})$ (c.f. (3.6)), will converge exponentially to constants that depend only on the edges $\gamma_{\pm}$ of the support of $\varrho_{\mathbf{b}}$, as shown by Theorem 3.1. This, combined with the perturbation analysis of [15], shows that the $n\times n$ principal submatrices of $\mathcal{L}(\widehat{\mu}_{\mathbf{b}})$ and $\mathcal{L}(\mu_{W,\mathbf{b}})$ are relatively close for $n\ll N^{1/6}$, as detailed in Corollary 5.3. As a result, $\mu_{W,\mathbf{b}}$ exhibits simple asymptotic properties that allow for efficient recovery of information about the ASD and the outliers of $W$ through the spiked ASD.
@@
Given the strong asymptotic relationship between $\mu_{W,\mathbf{b}}$ and $\widehat{\mu}_{\mathbf{b}}$, the Lanczos algorithm is applied until a steady state is reached, at which point the matrix $J_{n}$ is extended by constants to be semi-infinite (we refer the reader to $\mathcal{J}_{0}=\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ in the proof of Theorem 5.5). From this semi-infinite Jacobi matrix $\mathcal{J}_{0}$, we define the spectral measure $\widehat{\mu}_{0}$ as [10],
@@
[…]
@@
Let $\widehat{m}_{0}(z)$ be the estimate after running our algorithm (1.14) for $n=\mathrm{O}(\log N)$ iterations. Moreover, let $\widehat{\gamma}_{\pm}$ be estimates of the support endpoints and $\widehat{\gamma}_{j}$ ($j=1,2,\dots,\widehat{r}$) be the poles of $\widehat{m}_{0}(z)$ for $z>\widehat{\gamma}_{+}+N^{-1/\delta}$ with $20<\delta<1/$ 2, then the true number of spikes $r$ is equal to $\widehat{r}$, with overwhelming probability.
@@
  • Detection based on the first few outlier eigenvalues. It is well known that if spikes exist, the corresponding outlier eigenvalues may undergo the so-called BBP transition [4, 13]. In particular, if the spikes (i.e., $\widetilde{\sigma}_{i}$ in (1.5)) exceed a certain threshold, the corresponding eigenvalues $\lambda_{i}$ in (1.6) will separate from the support of the ASD. This suggests that under appropriate assumptions on the spikes (see Remark 4.5 below), one can estimate the number of spikes by counting the number of outlier eigenvalues departing from the ASD. Several estimators have been developed based on this idea, depending on how the outliers are counted. For instance, [5, 17, 18, 37] used the ratio of differences between consecutive eigenvalues, while [12, 41, 40, 49] employed eigen-ratios or eigen-differences. Additionally, [31] applied a nonparametric approach to identify the separation gap between outliers and bulk eigenvalues, and [16] utilized a bootstrap procedure. A common challenge in these methods is the need to carefully select a tuning threshold to distinguish larger outlier eigenvalues from the smaller, more rigid bulk eigenvalues. This threshold is often determined through Monte Carlo simulations and sometimes requires the knowledge of $\Sigma_{0}$ in (1.9), which can be computationally expensive and lack robustness.
@@
  • Procedures based on all or nearly all the eigenvalues. In this category of approaches, instead of directly counting the larger outlier eigenvalues, which requires a precise threshold, methods typically rely on utilizing nearly all eigenvalues. For example, [2, 36] apply information criteria in model selection (such as AIC and BIC), making use of all eigenvalues; [20] employs a parallel analysis approach that also utilizes all eigenvalues; and [29] reconstructs the limiting spectrum based on eigenvalue rigidity, using nearly all eigenvalues. Additionally, there is a research direction focused on estimating the number of spikes via hypothesis testing, which depends on all eigenvalues in various forms of linear spectral statistics (LSS); see, for example, [6, 19, 25, 28, 38]. Many of these methods require prior knowledge of the entry distribution or the structure of $\Sigma_{0}.$ Moreover, since these approaches rely on computing nearly all eigenvalues, their computational complexity increases significantly as the sample size and dimension grows.
@@
As discussed and compared in [29], the second category of methods, which utilize all or nearly all eigenvalues, is generally the most numerically robust and relies on weaker eigen-gap assumptions compared to the first category, which only uses a few outlier eigenvalues. However, from a computational perspective, methods in the first category—those relying on the largest few eigenvalues—requires fewer calculations, typically incurring a cost of order $\mathrm{O}(bcN^{2}\log N)$. Here, $b$ is a potentially diverging constant that depends on the eigen-gap of the extreme eigenvalues, while $c$ is usually divergent and represents the complexity of selecting the threshold in Monte Carlo simulations. In contrast, methods in the second category, which use all or nearly all eigenvalues, require at least a cost of order $\mathrm{O}(N^{3})$ to compute the spectrum of $W$.
@@
The structure of the paper is as follows. In Section 2, we review the well-known Lanczos algorithm and introduce our proposed algorithms. Section 3 discusses the details of orthogonal polynomials associated with a given measure and their connection to the corresponding Jacobi matrices and their Cholesky factorizations. Additionally, we provide the fixed-point equation for characterizing the Stieltjes transform via the Cholesky factorization. In Section 4, we formalize our model and present the relevant random matrix theory needed to analyze the VESD, which is central to our theoretical developments. Sections 5 and 6 focus on the main analysis of our proposed algorithms. Finally, in Section 7, we demonstrate the method and present further numerical investigations. The code used to generate the plots in this paper is available at [1].
@@ 2.1. The Lanczos algorithm
The Lanczos algorithm (c.f., [34, 47]) is an iterative method used to obtain a tridiagonal approximation of a symmetric or Hermitian matrix through matrix-vector multiplications and inner products. In its simplest form, it is given by Algorithm B.1 in Appendix A.1. In our computations we use Lanczos with reorthogonalization [39].
@@
where $(\lambda_{j})_{j=1}^{n}$ are the eigenvalues of $J_{n}$ and $\omega_{j}$ is the squared modulus of the first component of the normalized eigenvector associated to $\lambda_{j}$. The spectral measure $\mu_{J}$, with $n=N$ and $J=J_{N}(W,\mathbf{b})$, coincides with the VESD associated with $W$ and $\mathbf{b}$ whenever $\mathbf{b}$ is a unit vector. There is a bijection between such measures and Jacobi matrices [11], making the Lanczos algorithm an effective computational method for iteratively determining the Jacobi matrix entries associated with $\mu_{W,\mathbf{b}}$ without requiring the full spectrum of $W$.
@@
(2.2) $L_{n}:=\begin{bmatrix}\alpha_{0}\\ \beta_{0}&\alpha_{1}\\ &\ddots&\ddots\\ &&\beta_{n-1}&\alpha_{n-1}\end{bmatrix}.$

### v2

@@ 1. Introduction
Large sample covariance matrices play a fundamental role in high-dimensional data analysis. A widely studied and sophisticated model in the literature is the spiked covariance matrix model [Ding2021c, Johnstone2001], in which a finite number of spikes—eigenvalues that separate from the bulk of the spectrum—are introduced into the population covariance matrix $\Sigma$. In this model, a key task is to estimate the number of spikes using the associated sample covariance matrices. Such a problem has been studied in various settings; see Section 1.3 for a detailed review. However, all existing works leverage the spectral properties of $W$, relying on specific statistics constructed from its eigenvalues—either the prominent outlier eigenvalues or the relatively small bulk eigenvalues. Despite the potential statistical consistency of these approaches under certain conditions, they can be computationally intensive and prone to numerical errors due to the direct calculation of eigenvalues [^1] Modern eigensolvers are very accurate for symmetric matrices. The errors here stem from using individual eigevalues that may fluctuate wildly, while averages of eigenvalues have much smaller variance. and associated Monte Carlo estimations.
@@
In this paper, instead of relying on eigenvalue-based statistics, we introduce a novel algorithmic approach built upon the well-known Lanczos iteration [^2] Due to instabilities in the iteration we use reorthogonalization [Paige1980]. [Lanczos1950], which operates directly on the sample covariance matrix $W$ using random vectors, uniformly distributed on the hypersphere. Our proposed methods are both statistically consistent and numerically cheap while still being robust. Furthermore, as a byproduct of determining the number of spikes, we also obtain estimates for the asymptotic density of the empirical spectral distribution of $W$ and the asymptotic locations of the outliers in $W$.
@@
Our approach is based on the observation that, under mild assumptions, the number of spikes is, with high probability, identical to the number of poles of a complex function—specifically, the Stieltjes transform of a measure that connects the empirical spectral distribution (ESD) and its asymptotic counterpart, the asymptotic spectral distribution (ASD). This stochastic measure, which we refer to as the spiked ASD (see (1.11) and (5.1) below), preserves a scaled version of the continuous density of the ASD while accounting for the random outliers in the ESD. Consequently, a key challenge is to develop an efficient and robust estimator for the spiked ASD. Rather than directly relying on the eigenvalues of $W$, the central innovation of our method is the use of Cholesky factorization for the Jacobi matrix obtained by applying the Lanczos iteration with $W$ to several test vectors. This approach is motivated by the fact that the Stieltjes transform of the spiked ASD of $W$ can be approximately characterized by a new fixed-point equation, based on continued fraction theory, whose coefficients are derived from the entries of the Cholesky factorization of the Jacobi matrix associated with the spiked ASD (see Section 3.2). This fixed-point equation differs from the commonly-used self-consistent equation in random matrix theory [Bai2010, Knowles2017]. Crucially, this equation, along with the Lanczos algorithm, enables a numerical estimation of the Stieltjes transform of the ASD through an iterative process without directly computing the eigenvalues of $W$ (see Algorithm SR.2 and Section 3.2).
@@
On the techinical side, our proofs rely on three main ideas. The first one is the fact that the orthogonal polynomials for a generic limiting random matrix spectral distribution exhibit exponentially small error terms in their asymptotics [Kuijlaars2003, Yattselev2015, Ding2021b]. These results are based on the analysis of the Fokas–Its–Kitaev Riemann–Hilbert characterization of orthogonal polynomials [FokasItsKitaev]. The second is the so-called anisotropic local laws for covariance matrices [Knowles2017, Pillai2014] and their extension to spiked models [Ding2019]. The last technique is a recently developed perturbation theory [Ding2021b] for orthogonal polynomials, which builds on the Fokas–Its–Kitaev Riemann–Hilbert theory and is particularly compatible with the local laws from random matrix theory.
@@
Figure 1. Left: The ESD of $\Sigma$, where $\Sigma$ is an $N\times N$ diagonal matrix with $N=6000$ and entries given by the quantiles of the density $\frac{1}{K}\frac{x^{4}+1}{x^{2}}\sqrt{x-0.1}\sqrt{4-x}$ with $K$ being a normalizing constant. The first three diagonal entries are modified to 7, 6, and 6, forming the spikes of $\Sigma$. Right: The ESD of $W$ defined in (1.1), with $c_{N}=0.1$, $M=60000$, and $X$ having iid standard normal entries. The ESD is compared against the estimated outliers, their locations, and the approximate ASD obtained using Algorithms P.2 and P.3 with parameters $k=150$ and $C=1$.
@@ 1.1. The model and problem formulation
where $\Sigma$ is a positive-definite deterministic matrix, referred to as the population covariance matrix, and $X$ is an $N\times M$ random matrix with centered, independent, and identically distributed (iid) entries. In this paper, we are interested in the high-dimensional setting, where $M$ is comparably large to $N$. Specifically, we assume the existence of a small constant $0<\tau<1$ such that the aspect ratio $c_{N}:=N/M$ satisfies
@@
We adopt the spiked covariance matrix model following the framework established in [Ding2021]. Let the spiked covariance matrix $\Sigma$ admit the following spectral decomposition
@@
We differentiate $\Sigma_{0}$ from $\Sigma$ because if a limit is desired for certain spectral statistics of (1.1), then additional assumptions will need to be imposed on $\Sigma_{0}$; see Assumption 2 for more details. As a historical note, we recall that if $\sigma_{i}\equiv 1$ for $1\leq i\leq N$, it is well known that the eigenvalues of $W_{0}$ obey the Marchenko-Pastur law [Marcenko1967]. For general $\Sigma_{0}$, they are governed by the so-called deformed Marchenko-Pastur law [Bai2010, Knowles2017], which, in turn, determines $\mu_{\mathrm{ASD}}$.
@@ 1.2. An overview of our approach
The entries of the semi-infinite Jacobi matrix $\mathcal{J}(\widehat{\mu}_{\mathbf{b}})$ associated with $\widehat{\mu}_{\mathbf{b}}$ (see (3.4)) and its Cholesky factor $\mathcal{L}(\widehat{\mu}_{\mathbf{b}})$ (c.f. (3.6)), will converge exponentially to constants that depend only on the edges $\gamma_{\pm}$ of the support of $\varrho_{\mathbf{b}}$, as shown by Theorem 3.1. This, combined with the perturbation analysis of [Ding2021b], shows that the $n\times n$ principal submatrices of $\mathcal{L}(\widehat{\mu}_{\mathbf{b}})$ and $\mathcal{L}(\mu_{W,\mathbf{b}})$ are relatively close for $n\ll N^{1/6}$, as detailed in Corollary 5.3. As a result, $\mu_{W,\mathbf{b}}$ exhibits simple asymptotic properties that allow for efficient recovery of information about the ASD and the outliers of $W$ through the spiked ASD.
@@
Given the strong asymptotic relationship between $\mu_{W,\mathbf{b}}$ and $\widehat{\mu}_{\mathbf{b}}$, the Lanczos algorithm is applied until a steady state is reached, at which point the matrix $J_{n}$ is extended by constants to be semi-infinite (we refer the reader to $\mathcal{J}_{0}=\mathcal{L}_{0}\mathcal{L}_{0}^{*}$ in the proof of Theorem 5.5). From this semi-infinite Jacobi matrix $\mathcal{J}_{0}$, we define the spectral measure $\widehat{\mu}_{0}$ as [DeiftOrthogonalPolynomials],
@@
(1.14) $(W,\mathbf{b}\sim\mathrm{Unif}(\mathbb{S}^{N-1}))\overset{\begin{subarray}{c}\text{Lanczos with}\\ n~\text{steps}\end{subarray}}{\longrightarrow}J_{n}\overset{\text{extend}}{\longrightarrow}\mathcal{J}_{0}=\mathcal{L}_{0}\mathcal{L}_{0}^{*}\overset{\begin{subarray}{c}\text{continued}\\ \text{fractions}\end{subarray}}{\longrightarrow}\widehat{m}_{0}\longrightarrow\widehat{\mu}_{0}\approx\widehat{\mu}_{\mathbf{b}}\approx\mu_{\mathrm{sASD}}.$
@@
Let $\widehat{m}_{0}(z)$ be the estimate after running our algorithm (1.14) for $n=\mathrm{O}(\log N)$ iterations. Moreover, let $\widehat{\gamma}_{\pm}$ be estimates of the support endpoints and $\widehat{\gamma}_{j}$ ($j=1,2,\dots,\widehat{r}$) be the poles of $\widehat{m}_{0}(z)$ for $z>\widehat{\gamma}_{+}+N^{-1/\delta}$ with $0<\delta<1/2$, then the true number of spikes $r$ is equal to $\widehat{r}$, with overwhelming probability.
@@
  • Detection based on the first few outlier eigenvalues. It is well known that if spikes exist, the corresponding outlier eigenvalues may undergo the so-called BBP transition [Baik2005, Ding2021c]. In particular, if the spikes (i.e., $\widetilde{\sigma}_{i}$ in (1.5)) exceed a certain threshold, the corresponding eigenvalues $\lambda_{i}$ in (1.6) will separate from the support of the ASD. This suggests that under appropriate assumptions on the spikes (see Remark 4.5 below), one can estimate the number of spikes by counting the number of outlier eigenvalues departing from the ASD. Several estimators have been developed based on this idea, depending on how the outliers are counted. For instance, [Bao2015Universality, Ding2019, Ding2022TracyWidom, onatski2009testing] used the ratio of differences between consecutive eigenvalues, while [10.3150/19-BEJ1129, passemier2012determining, passemier2014estimation, Vinogradova2013StatisticalInference] employed eigen-ratios or eigen-differences. Additionally, [kritchman2009nonparametric] applied a nonparametric approach to identify the separation gap between outliers and bulk eigenvalues, and [dingboostrap] utilized a bootstrap procedure. A common challenge in these methods is the need to carefully select a tuning threshold to distinguish larger outlier eigenvalues from the smaller, more rigid bulk eigenvalues. This threshold is often determined through Monte Carlo simulations and sometimes requires the knowledge of $\Sigma_{0}$ in (1.9), which can be computationally expensive and lack robustness.
@@
  • Procedures based on all or nearly all the eigenvalues. In this category of approaches, instead of directly counting the larger outlier eigenvalues, which requires a precise threshold, methods typically rely on utilizing nearly all eigenvalues. For example, [Bai2018ConsistencyOA, Nadakuditi2008Sample] apply information criteria in model selection (such as AIC and BIC), making use of all eigenvalues; [PA] employs a parallel analysis approach that also utilizes all eigenvalues; and [Ke2023] reconstructs the limiting spectrum based on eigenvalue rigidity, using nearly all eigenvalues. Additionally, there is a research direction focused on estimating the number of spikes via hypothesis testing, which depends on all eigenvalues in various forms of linear spectral statistics (LSS); see, for example, [Bianchi2011Performance, Dobriban2017SharpDI, 10551866, johnstone2018testing, onatski2014signal]. Many of these methods require prior knowledge of the entry distribution or the structure of $\Sigma_{0}.$ Moreover, since these approaches rely on computing nearly all eigenvalues, their computational complexity increases significantly as the sample size and dimension grows.
@@
As discussed and compared in [Ke2023], the second category of methods, which utilize all or nearly all eigenvalues, is generally the most numerically robust and relies on weaker eigen-gap assumptions compared to the first category, which only uses a few outlier eigenvalues. However, from a computational perspective, methods in the first category—those relying on the largest few eigenvalues—requires fewer calculations, typically incurring a cost of order $\mathrm{O}(bcN^{2}\log N)$. Here, $b$ is a potentially diverging constant that depends on the eigen-gap of the extreme eigenvalues, while $c$ is usually divergent and represents the complexity of selecting the threshold in Monte Carlo simulations. In contrast, methods in the second category, which use all or nearly all eigenvalues, require at least a cost of order $\mathrm{O}(N^{3})$ to compute the spectrum of $W$.
@@
The structure of the paper is as follows. In Section 2, we review the well-known Lanczos algorithm and introduce our proposed algorithms. Section 3 discusses the details of orthogonal polynomials associated with a given measure and their connection to the corresponding Jacobi matrices and their Cholesky factorizations. Additionally, we provide the fixed-point equation for characterizing the Stieltjes transform via the Cholesky factorization. In Section 4, we formalize our model and present the relevant random matrix theory needed to analyze the VESD, which is central to our theoretical developments. Sections 5 and 6 focus on the main analysis of our proposed algorithms. Finally, in Section 7, we demonstrate the method and present further numerical investigations. The code used to generate the plots in this paper is available at [AbiYounes2025SpikeDetection].
@@ 2.1. The Lanczos algorithm
The Lanczos algorithm (c.f., [Lanczos1950, TrefethenBau]) is an iterative method used to obtain a tridiagonal approximation of a symmetric or Hermitian matrix through matrix-vector multiplications and inner products. In its simplest form, it is given by Algorithm B.1 in Appendix A.1. In our computations we use Lanczos with reorthogonalization [Paige1980].
@@
where $(\lambda_{j})_{j=1}^{n}$ are the eigenvalues of $J_{n}$ and $\omega_{j}$ is the squared modulus of the first component of the normalized eigenvector associated to $\lambda_{j}$. The spectral measure $\mu_{J}$, with $n=N$ and $J=J_{N}(W,\mathbf{b})$, coincides with the VESD associated with $W$ and $\mathbf{b}$ whenever $\mathbf{b}$ is a unit vector. There is a bijection between such measures and Jacobi matrices [DeiftEigenvalue], making the Lanczos algorithm an effective computational method for iteratively determining the Jacobi matrix entries associated with $\mu_{W,\mathbf{b}}$ without requiring the full spectrum of $W$.
@@
(2.2) $L_{n}:=\begin{bmatrix}\alpha_{0}\\ \beta_{0}&\alpha_{1}\\ &\ddots&\ddots\\ &&\beta_{n-2}&\alpha_{n-1}\end{bmatrix}$.